OpenStudy (anonymous):

Solve the equation (y+2/y)^2 + 3(y+2/y) = 4 using substitution method.

OpenStudy (eyust707):

hmm do you know what the substitution method is?

OpenStudy (anonymous):

yea

OpenStudy (eyust707):

what is it i honestly forgot..

OpenStudy (eyust707):

hmm i was reading that its when you have 2 variables, solve for one, then sub tht one into the other..

OpenStudy (anonymous):

yea i was gettin u an exact definition

OpenStudy (eyust707):

ok ok

OpenStudy (eyust707):

well we could say that (y+2/y) ^2 = 1

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

if you notice that would solve it.

OpenStudy (eyust707):

actually no nvm

OpenStudy (eyust707):

thats not right... lol i can solve it how i normally would but not sure what they mean by substiution

OpenStudy (eyust707):

ohh maybe they want you to say x= (y + 2) / y and that x^2 + 3x = 4

OpenStudy (eyust707):

x^2 +3x -4 = 0

OpenStudy (anonymous):

ughh no clue

OpenStudy (eyust707):

(x +4) (x - 1) = 0

OpenStudy (eyust707):

so x = -4 or x = 1

OpenStudy (eyust707):

well whats x?

OpenStudy (eyust707):

we said that in the begining

OpenStudy (eyust707):

x= (y + 2) / y

OpenStudy (eyust707):

(y + 2) / y = 4 and (y + 2) / y = -1

OpenStudy (eyust707):

solve for y

OpenStudy (eyust707):

that used substituion cause i subbed in y for x =)

OpenStudy (anonymous):

wait which one is right

OpenStudy (eyust707):

there are two possible answers tht satifiy the equation

OpenStudy (anonymous):

ok thanks lol i suck at math

OpenStudy (eyust707):

its coo lol if you want call me on skype we could prolly mob thru way more problems.. ill prolly be on for the next hour or so

OpenStudy (anonymous):

yay cus i got like 4 problems left...wats ur skype name

OpenStudy (eyust707):

its the same as on here eyust707

OpenStudy (anonymous):

ok sent it

OpenStudy (eyust707):

call mee

OpenStudy (anonymous):

ok give me one min

OpenStudy (eyust707):

kk